# Motion Polynomials Admitting a Factorization with Linear Factors

### v3

@@ Abstract
Zijia Li11footnotemark: 1 Hans-Peter Schröcker22footnotemark: 2 Mikhail Skopenkov33footnotemark: 3 Daniel F. Scharler22footnotemark: 2, [^4] Daniel F. Scharler greatly contributed to an early version of this manuscript. He tragically passed away in April 2022 at the age of only 29.
@@ 1 Introduction
The first publication on motion polynomial factorization is [9]. There, the authors showed that a “generic” motion polynomial could be written in finitely many ways as a product of monic linear motion polynomials. They also related these factorizations to a decomposition of rational motions into coupled rotations and converted it into the construction of mechanical linkages. For instance, using different factorizations of a cubic motion polynomial, the authors in [9] found a new family of paradoxically mobile closed-loop linkages with six revolute axes. Notice that the family contained the well-known Bricard Octahedra of type three [7] as a sub-class. Computational design issues were the topic of [10]. In [6, 20], the factorization of bounded motion polynomials was used for the construction of rational Kempe linkages. The potential for applications in architectural design is studied in [15]. Moreover, in [24], it was used to synthesize single-loop variable-degree-of-freedom mechanisms.
@@ 2 Preliminaries
We denote by $\operatorname{realgcd}(A,B)$ the greatest common real monic polynomial divisor of $A$ and $B$. In addition, we define $\operatorname{realgcd}(M):=\operatorname{realgcd}(M,M)$ and $\operatorname{realgcd}(0):=1$. In case of $\operatorname{realgcd}(M)=1$, we say that $M$ is reduced.
@@
A monic reduced motion polynomial $M=P+\varepsilon D$ is generic if $\operatorname{realgcd}(P)=1$. It is called bounded if $\operatorname{realgcd}(P)$ has no real zero, and unbounded otherwise.
@@
If a monic polynomial $F\in\mathbb{R}[t]$ divides the product $AB$ of polynomials $A,B\in\mathbb{H}[t]$ but $\operatorname{realgcd}(F,A)=\operatorname{realgcd}(F,B)=1$, then $\operatorname{rgcd}(F,A)=\operatorname{lgcd}(F,B)^{*}$ and $F=\operatorname{rgcd}(F,A)\operatorname{lgcd}(F,B)$.
@@
The polynomial $F$ has no real roots, otherwise, it would have a common real root with $A$ or $B$, contradicting $\operatorname{realgcd}(F,A)=\operatorname{realgcd}(F,B)=1$. Let $N$ be a real irreducible quadratic factor of $F$. Since $N\,|\,F\,|\,AB$, it follows that $N\,|\,A{A}^{\ast}B{B}^{\ast}$, hence $N\,|\,A{A}^{\ast}$ or $N\,|\,B{B}^{\ast}$. Assume that $N\,|\,A{A}^{\ast}$ without loss of generality. Using division with remainders, we find $Q,R,B_{1}\in\mathbb{H}[t]$ and $r\in\mathbb{H}$ such that $A=QN+R$ and $B={R}^{\ast}B_{1}+r$. Here $R\neq 0$ because $\operatorname{realgcd}(F,A)=1$. Since $N\,|\,A{A}^{\ast}$, it follows that $R{R}^{\ast}=pN$ for some nonzero $p\in\mathbb{R}$. Since $N\,|\,AB$, it follows that $N\,|\,Rr$, hence $r=0$. The polynomials $A_{1}:=AR^{-1}=Q{R}^{\ast}/p+1$, $B_{1}=({R}^{\ast})^{-1}B$, and $F_{1}:=F/N$ still satisfy the assumptions of the lemma. Applying the same construction inductively, we get $A_{n},B_{n},R_{n}\in\mathbb{H}[t]$ such that $A=A_{n}R_{n}$, $B={R_{n}}^{\ast}B_{n}$, $F=R_{n}{R_{n}}^{\ast}$, and $R_{n}$ is monic.
It remains to show that $R_{n}=\operatorname{rgcd}(F,A)$ and ${R_{n}}^{\ast}=\operatorname{lgcd}(F,B)$ (cf. [18, Lemma 3]). Indeed, by the extended Euclidean algorithm, $\operatorname{rgcd}(F,A)=XA+YF$ for some $X,Y\in\mathbb{H}[t]$. Hence $\operatorname{rgcd}(F,A)=SR_{n}$ for some $S\in\mathbb{H}[t]$. Since $F={R_{n}}^{\ast}R_{n}$ and $A=A_{n}R_{n}$ are right-divisible by $\operatorname{rgcd}(F,A)=SR_{n}$, it follows that ${R_{n}}^{\ast}$ and $A_{n}$ are right-divisible by $S$. Then $R_{n}$ is left-divisible by ${S}^{\ast}$, hence $F$ and $A$ are divisible by $S{S}^{\ast}$. Since $\operatorname{realgcd}(F,A)=1$, it follows that $S=1$ and $R_{n}=\operatorname{rgcd}(F,A)$. Analogously, ${R_{n}}^{\ast}=\operatorname{lgcd}(F,B)$. ∎
@@ 3.1 Factorizations of Unbounded Motion Polynomials
Assume $\operatorname{realgcd}(P)$ has a linear real factor $p$ of multiplicity two. We can factor $M$ into $M_{1}M_{2}$, where $M_{1}=pP_{1}+\varepsilon D_{1}$ and $M_{2}=pP_{2}+\varepsilon D_{2}$ by expanding the product of linear factors into two parts. But the product $M_{1}M_{2}=p^{2}P_{1}P_{2}+\varepsilon p(D_{1}P_{2}+P_{1}D_{2})$ is not reduced. We have a contradiction. ∎
In what follows, we consider only bounded motion polynomials $M=P+\varepsilon D$, where $P=cQ$ with monic $c\in\mathbb{R}[t]$ and $Q\in\mathbb{H}[t]$ such that $\operatorname{realgcd}(Q)=1$ and $c$ has no real zeros.
@@ 3.2 Factorizations of Generic Motion Polynomials
8:$M\leftarrow\operatorname{rquo}(M,t-h)$
@@
Let $M=P+\varepsilon D\in\mathbb{DH}[t]$ be a monic motion polynomial, $G\in\mathbb{R}[t]$ be a monic polynomial such that $G|\nu(M)$ and $\operatorname{realgcd}(G,P)=1$. Then there is a motion polynomial $M_{1}$ that admits a factorization into monic linear factors such that $G=M_{1}{M_{1}}^{\ast}$ and $M=M_{1}M_{2}$ for some motion polynomial $M_{2}$.
@@
The polynomial $G$ has no real roots because $G|\nu(P)$ and $\operatorname{realgcd}(G,P)=1$. The linear factors are obtained using Algorithm 1 by the recursive procedure where each quadratic real factor of $G$ gives a linear factor. ∎
Beware that if a non-generic motion polynomial $M$ has a decomposition into monic linear motion polynomial factors, $M_{2}$ need not have one.
@@ 3.3 Decomposition of Bounded Motion Polynomials into Factors with Primary Norm
  • There exists a polynomial $S\in\mathbb{R}[t]$ of degree $\deg S\leq\deg\operatorname{realgcd}(P)$ such that $MS$ admits a factorization with linear motion polynomial factors.
@@
7:$Q\leftarrow\frac{P}{\operatorname{realgcd}(P)}$, $c_{2}\leftarrow\operatorname{realgcd}(P,N^{n})$, and $c_{1}\leftarrow\operatorname{realgcd}(P)/c_{2}$
8:$Q_{2}\leftarrow\operatorname{rgcd}(N^{n}/c_{2}^{2},Q)$ and $Q_{1}\leftarrow\operatorname{rquo}(Q,Q_{2})$
@@
$\nu(M^{\prime})$ $=\left(c_{1}Q_{1}+\varepsilon\frac{Q_{1}D_{1}}{\nu(D_{1})}\right){\left(c_{1}Q_{1}+\varepsilon\frac{Q_{1}D_{1}}{\nu(D_{1})}\right)}^{\ast}=c_{1}^{2}\nu(Q_{1})+\varepsilon c_{1}\nu(Q_{1})\underbrace{(D_{1}+{D}^{\ast}_{1})}_{=0}=\prod_{i=1}^{r-1}N_{i}^{n_{i}},$
$\nu(M_{r})$ $=\left(c_{2}Q_{2}+\varepsilon\frac{D_{2}Q_{2}}{\nu(Q_{2})}\right){\left(c_{2}Q_{2}+\varepsilon\frac{D_{2}Q_{2}}{\nu(Q_{2})}\right)}^{\ast}=c_{2}^{2}\nu(Q_{2})+\varepsilon c_{2}\nu(Q_{2})\underbrace{(D_{2}+{D}^{\ast}_{2})}_{=0}=N_{r}^{n_{r}},$
@@ 3.4 Factorizations of Bounded Motion Polynomials with Primary Norm
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with monic $c\in\mathbb{R}[t]$, $\operatorname{realgcd}(Q)=1$, and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. If $M$ admits a factorization into monic linear motion polynomial factors, then there exist quaternion polynomials $Q_{L}$, $Q_{R}$, $D_{L}$, $D_{R}$, $D^{\prime}\in\mathbb{H}[t]$ such that $M=(Q_{L}+\varepsilon D_{L})(c+\varepsilon D^{\prime})(Q_{R}+\varepsilon D_{R})$, where $Q_{L}+\varepsilon D_{L}$, $c+\varepsilon D^{\prime}$, and $Q_{R}+\varepsilon D_{R}$ are products of monic linear motion polynomials.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $\operatorname{realgcd}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. Let $M$ admit a factorization into monic linear motion polynomial factors and consider polynomials $Q_{L}$ and $Q_{R}$ from Lemma 3. Define $g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ and $g_{R}\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast})\in\mathbb{R}[t]$. Then $g_{L}$ divides $\nu(Q_{L})$ and $g_{R}$ divides $\nu(Q_{R})$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $\operatorname{realgcd}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. If $M$ admits a factorization into monic linear motion polynomial factors, then $cg_{L}$ or $cg_{R}$ divides $\nu(D)$, where $g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ and $g_{R}\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast})$.
@@
1:A bounded monic reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$, with $\deg M=n\geq 1$, $c=\operatorname{realgcd}P$, $\deg(c)>0$, $P=cQ$, and $\nu(M)=N^{n}$ for a real irreducible quadratic polynomial $N$ such that $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ divides $\operatorname{realgcd}(c,D{Q}^{\ast})$ and $cg$ divides $\nu(D)$
@@
For a bounded, monic, reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$, with $\deg M=n>1$, $c=\operatorname{realgcd}P$, $\deg(c)>0$, $P=cQ$, and $\nu(M)=N^{n}$ for a real, irreducible, quadratic polynomial $N$ such that $g=g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ divides $\operatorname{realgcd}(c,D{Q}^{\ast})$ and $cg$ divides $\nu(D)$, Algorithm 4 outputs a triple $(M_{L},M_{C},M_{R})$ of monic motion polynomials such that $M=M_{L}M_{C}M_{R}$, where $M_{L}$ and $M_{R}$ are generic motion polynomials, $M_{C}=c+\varepsilon D^{\prime}$ is a product of two generic motion polynomials, and $D^{\prime}\in\mathbb{H}[t]$.
@@
From Steps 1-3, if $\deg(Q)=0$, by the assumption, then $M=c+\varepsilon D$ and we have $g=1$ and $c|\nu(D)$. Applying Lemma 2 to $A=D$, $B={D}^{\ast}$, and $F=c$, we conclude that $D=M_{1}M_{2}$ with $M_{1}=\operatorname{lgcd}(c,D)$ and some $M_{2}\in\mathbb{H}[t]$ so that $c=M_{1}{M}^{\ast}_{1}$. Then $M=M_{1}{M}^{\ast}_{1}+\varepsilon M_{1}M_{2}=M_{1}({M}^{\ast}_{1}+\varepsilon M_{2})$ is a product of two generic motion polynomials. Assume further $\deg(Q)\neq 0$.
@@
  • We claim that $c|D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}$. Since $c=g$, we have $c|Q{Q}^{\ast}|{D}^{\ast}_{L}Q{Q}^{\ast}D_{L}$ and $c|Q_{L}{Q}^{\ast}_{L}\frac{D{D}^{\ast}}{g^{2}}$. Since $D=Q_{L}D_{2}$ and $g^{2}|{D}^{\ast}D$, it follows that $g|D_{2}{D}^{\ast}_{2}$. Since $M$ is reduced, it follows that $\operatorname{realgcd}(g,D_{2})=\operatorname{realgcd}(g,D)=1$, and $g=\nu(\operatorname{rgcd}(D_{2},g))=\nu(\operatorname{rgcd}(D,g))$ by Lemma 2, hence $\operatorname{rgcd}(D_{2},g)=\operatorname{rgcd}(D,g)$. Since $g|D{Q}^{\ast}$, it follows that $g=\operatorname{rgcd}(g,D)\operatorname{lgcd}(g,{Q}^{\ast})=\operatorname{rgcd}(g,D_{2})\operatorname{lgcd}(g,{Q}^{\ast})$ by Lemma 2, thus $g|D_{2}{Q}^{\ast}$. Therefore, $g^{2}$ divides $Q{D}^{\ast}Q_{L}$ and ${Q}^{\ast}_{L}{D}{Q}^{\ast}$. Therefore, $c$ divides $D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}$ because we have $D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}={D}^{\ast}_{L}Q{Q}^{\ast}D_{L}+Q_{L}{Q}^{\ast}_{L}\frac{D{D}^{\ast}}{g^{2}}+\frac{c}{g^{2}}\left({D}^{\ast}_{L}Q{D}^{\ast}Q_{L}+{Q}^{\ast}_{L}{D}{Q}^{\ast}D_{L}\right)$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n\in\mathbb{N}$ with $\operatorname{realgcd}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$. If $cg$ divides $\nu(D)$, then $M$ admits a factorization into monic linear factors.

### v4

@@ Abstract
Zijia Li11footnotemark: 1 Hans-Peter Schröcker22footnotemark: 2 Mikhail Skopenkov33footnotemark: 3 Daniel F. Scharler22footnotemark: 2, [^4] Daniel F. Scharler greatly contributed to an earlier version of this manuscript. He tragically passed away on April 12, 2022 at the age of only 29.
@@ 1 Introduction
The first publication on motion polynomial factorization is [9]. There, the authors showed that a “generic” monic motion polynomial could be written in finitely many ways as a product of monic linear motion polynomials. They also related these factorizations to a decomposition of rational motions into coupled rotations and converted them into the construction of mechanical linkages. For instance, using different factorizations of a cubic motion polynomial, the authors in [9] found a new family of paradoxically mobile closed-loop linkages with six revolute axes. Notice that the family contained the well-known Bricard Octahedra of type three [7] as a sub-class. Computational design issues were the topic of [10]. In [6, 20], the factorization of bounded motion polynomials was used for the construction of rational Kempe linkages. The potential for applications in architectural design is studied in [15]. Moreover, in [24], it was used to synthesize single-loop variable-degree-of-freedom mechanisms.
@@ 2 Preliminaries
We denote by $\operatorname{mrpf}(A,B)$ the greatest common real monic polynomial divisor of $A$ and $B$. In addition, we define $\operatorname{mrpf}(M):=\operatorname{mrpf}(M,M)$ and $\operatorname{mrpf}(0):=1$. In case of $\operatorname{mrpf}(M)=1$, we say that $M$ is reduced.
@@
A monic reduced motion polynomial $M=P+\varepsilon D$ is generic if $\operatorname{mrpf}(P)=1$. It is called bounded if $\operatorname{mrpf}(P)$ has no real zero, and unbounded otherwise.
@@
If a monic polynomial $F\in\mathbb{R}[t]$ divides the product $AB$ of polynomials $A,B\in\mathbb{H}[t]$ but $\operatorname{mrpf}(F,A)=\operatorname{mrpf}(F,B)=1$, then $\operatorname{rgcd}(F,A)=\operatorname{lgcd}(F,B)^{*}$ and $F=\operatorname{rgcd}(F,A)\operatorname{lgcd}(F,B)$.
@@
The polynomial $F$ has no real roots, otherwise, it would have a common real root with $A$ or $B$, contradicting $\operatorname{mrpf}(F,A)=\operatorname{mrpf}(F,B)=1$. Let $N$ be a real irreducible quadratic factor of $F$. Since $N\,|\,F\,|\,AB$, it follows that $N\,|\,A{A}^{\ast}B{B}^{\ast}$, hence $N\,|\,A{A}^{\ast}$ or $N\,|\,B{B}^{\ast}$. Assume that $N\,|\,A{A}^{\ast}$ without loss of generality. Using division with remainders, we find $Q,R,B_{1}\in\mathbb{H}[t]$ and $r\in\mathbb{H}$ such that $A=QN+R$ and $B={R}^{\ast}B_{1}+r$. Here $R\neq 0$ because $\operatorname{mrpf}(F,A)=1$. Since $N\,|\,A{A}^{\ast}$, it follows that $R{R}^{\ast}=pN$ for some nonzero $p\in\mathbb{R}$. Since $N\,|\,AB$, it follows that $N\,|\,Rr$, hence $r=0$. The polynomials $A_{1}:=AR^{-1}=Q{R}^{\ast}/p+1$, $B_{1}=({R}^{\ast})^{-1}B$, and $F_{1}:=F/N$ still satisfy the assumptions of the lemma. Applying the same construction inductively, we get $A_{n},B_{n},R_{n}\in\mathbb{H}[t]$ such that $A=A_{n}R_{n}$, $B={R_{n}}^{\ast}B_{n}$, $F=R_{n}{R_{n}}^{\ast}$, and $R_{n}$ is monic.
It remains to show that $R_{n}=\operatorname{rgcd}(F,A)$ and ${R_{n}}^{\ast}=\operatorname{lgcd}(F,B)$ (cf. [18, Lemma 3]). Indeed, by the extended Euclidean algorithm, $\operatorname{rgcd}(F,A)=XA+YF$ for some $X,Y\in\mathbb{H}[t]$. Hence $\operatorname{rgcd}(F,A)=SR_{n}$ for some monic $S\in\mathbb{H}[t]$. Since $F={R_{n}}^{\ast}R_{n}$ and $A=A_{n}R_{n}$ are right-divisible by $\operatorname{rgcd}(F,A)=SR_{n}$, it follows that ${R_{n}}^{\ast}$ and $A_{n}$ are right-divisible by $S$. Then $R_{n}$ is left-divisible by ${S}^{\ast}$, hence $F$ and $A$ are divisible by $S{S}^{\ast}$. Since $\operatorname{mrpf}(F,A)=1$, it follows that $S=1$ and $R_{n}=\operatorname{rgcd}(F,A)$. Analogously, ${R_{n}}^{\ast}=\operatorname{lgcd}(F,B)$. ∎
@@ 3.1 Factorizations of Unbounded Motion Polynomials
Assume $\operatorname{mrpf}(P)$ has a linear real factor $p$ of multiplicity two. We can factor $M$ into $M_{1}M_{2}$, where $M_{1}=pP_{1}+\varepsilon D_{1}$ and $M_{2}=pP_{2}+\varepsilon D_{2}$ by expanding the product of linear factors into two parts. But the product $M_{1}M_{2}=p^{2}P_{1}P_{2}+\varepsilon p(D_{1}P_{2}+P_{1}D_{2})$ is not reduced. We have a contradiction. ∎
In what follows, we consider only bounded motion polynomials $M=P+\varepsilon D$, where $P=cQ$ with monic $c\in\mathbb{R}[t]$ and $Q\in\mathbb{H}[t]$ such that $\operatorname{mrpf}(Q)=1$ and $c$ has no real zeros.
@@ 3.2 Factorizations of Generic Motion Polynomials
8:$M\leftarrow\operatorname{lquo}(M,t-h)$
@@
Let $M=P+\varepsilon D\in\mathbb{DH}[t]$ be a monic motion polynomial, $G\in\mathbb{R}[t]$ be a monic polynomial such that $G|\nu(M)$ and $\operatorname{mrpf}(G,P)=1$. Then there is a motion polynomial $M_{2}$ that admits a factorization into monic linear factors such that $G=M_{2}{M_{2}}^{\ast}$ and $M=M_{1}M_{2}$ for some motion polynomial $M_{1}$.
@@
The polynomial $G$ has no real roots because $G|\nu(P)$ and $\operatorname{mrpf}(G,P)=1$. The linear factors are obtained using Algorithm 1 by the recursive procedure where each quadratic real factor of $G$ gives a linear factor. ∎
Beware that if a non-generic motion polynomial $M$ has a decomposition into monic linear motion polynomial factors, $M_{1}$ need not have one.
@@ 3.3 Decomposition of Bounded Motion Polynomials into Factors with Primary Norm
  • There exists a polynomial $S\in\mathbb{R}[t]$ of degree $\deg S\leq\deg\operatorname{mrpf}(P)$ such that $MS$ admits a factorization with linear motion polynomial factors.
@@
7:$Q\leftarrow\frac{P}{\operatorname{mrpf}(P)}$, $c_{2}\leftarrow\operatorname{realgcd}(P,N^{n})$, and $c_{1}\leftarrow\operatorname{mrpf}(P)/c_{2}$
8:$Q_{2}\leftarrow\operatorname{rgcd}(N^{n}/c_{2}^{2},Q)$ and $Q_{1}\leftarrow\operatorname{lquo}(Q,Q_{2})$
@@
$\nu(M^{\prime})$ $={\left(c_{1}Q_{1}+\varepsilon\frac{Q_{1}D_{1}}{\nu(Q_{1})}\right)}^{\ast}\left(c_{1}Q_{1}+\varepsilon\frac{Q_{1}D_{1}}{\nu(Q_{1})}\right)=c_{1}^{2}\nu(Q_{1})+\varepsilon c_{1}\underbrace{(D_{1}+{D}^{\ast}_{1})}_{=0}=\prod_{i=1}^{r-1}N_{i}^{n_{i}},$
$\nu(M_{r})$ $=\left(c_{2}Q_{2}+\varepsilon\frac{D_{2}Q_{2}}{\nu(Q_{2})}\right){\left(c_{2}Q_{2}+\varepsilon\frac{D_{2}Q_{2}}{\nu(Q_{2})}\right)}^{\ast}=c_{2}^{2}\nu(Q_{2})+\varepsilon c_{2}\underbrace{(D_{2}+{D}^{\ast}_{2})}_{=0}=N_{r}^{n_{r}},$
@@ 3.4 Factorizations of Bounded Motion Polynomials with Primary Norm
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with monic $c\in\mathbb{R}[t]$, $\operatorname{mrpf}(Q)=1$, and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. If $M$ admits a factorization into monic linear motion polynomial factors, then there exist quaternion polynomials $Q_{L}$, $Q_{R}$, $D_{L}$, $D_{R}$, $D^{\prime}\in\mathbb{H}[t]$ such that $M=(Q_{L}+\varepsilon D_{L})(c+\varepsilon D^{\prime})(Q_{R}+\varepsilon D_{R})$, where $Q_{L}+\varepsilon D_{L}$, $c+\varepsilon D^{\prime}$, and $Q_{R}+\varepsilon D_{R}$ are products of monic linear motion polynomials.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $\operatorname{mrpf}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. Let $M$ admit a factorization into monic linear motion polynomial factors and consider polynomials $Q_{L}$ and $Q_{R}$ from Lemma 3. Define $g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ and $g_{R}\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast})\in\mathbb{R}[t]$. Then $g_{L}$ divides $\nu(Q_{L})$ and $g_{R}$ divides $\nu(Q_{R})$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $\operatorname{mrpf}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. If $M$ admits a factorization into monic linear motion polynomial factors, then $cg_{L}$ or $cg_{R}$ divides $\nu(D)$, where $g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ and $g_{R}\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast})$.
@@
1:A bounded monic reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$, with $\deg M=n\geq 1$, $c=\operatorname{mrpf}P$, $\deg(c)>0$, $P=cQ$, and $\nu(M)=N^{n}$ for a real irreducible quadratic polynomial $N$ such that $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ divides $\operatorname{realgcd}(c,D{Q}^{\ast})$ and $cg$ divides $\nu(D)$
@@
For a bounded, monic, reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$, with $\deg M=n>1$, $c=\operatorname{mrpf}P$, $\deg(c)>0$, $P=cQ$, and $\nu(M)=N^{n}$ for a real, irreducible, quadratic polynomial $N$ such that $g=g_{L}\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D)$ divides $\operatorname{realgcd}(c,D{Q}^{\ast})$ and $cg$ divides $\nu(D)$, Algorithm 4 outputs a triple $(M_{L},M_{C},M_{R})$ of monic motion polynomials such that $M=M_{L}M_{C}M_{R}$, where $M_{L}$ and $M_{R}$ are generic motion polynomials, $M_{C}=c+\varepsilon D^{\prime}$ is a product of two generic motion polynomials, and $D^{\prime}\in\mathbb{H}[t]$.
@@
From Steps 1-3, if $\deg(Q)=0$, by the assumption, then $M=c+\varepsilon D$ and we have $g=1$ and $c|\nu(D)$. Applying Lemma 2 to $A={D}^{\ast}$, $B=D$, and $F=c$, we conclude that $D=M_{1}M_{2}$ with $M_{1}=\operatorname{lgcd}(c,D)$ and some $M_{2}\in\mathbb{H}[t]$ so that $c=M_{1}{M}^{\ast}_{1}$. Then $M=M_{1}{M}^{\ast}_{1}+\varepsilon M_{1}M_{2}=M_{1}({M}^{\ast}_{1}+\varepsilon M_{2})$ is a product of two generic motion polynomials. Assume further $\deg(Q)\neq 0$.
@@
  • We claim that $c|D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}$. Since $c=g$, we have $c|Q{Q}^{\ast}|{D}^{\ast}_{L}Q{Q}^{\ast}D_{L}$ and $c|Q_{L}{Q}^{\ast}_{L}\frac{D{D}^{\ast}}{g^{2}}$. Since $D=Q_{L}D_{2}$ and $g^{2}|{D}^{\ast}D$, it follows that $g|D_{2}{D}^{\ast}_{2}$. Since $M$ is reduced, it follows that $\operatorname{mrpf}(g,D_{2})=\operatorname{mrpf}(g,D)=1$, and $g=\nu(\operatorname{rgcd}(D_{2},g))=\nu(\operatorname{rgcd}(D,g))$ by Lemma 2, hence $\operatorname{rgcd}(D_{2},g)=\operatorname{rgcd}(D,g)$. Since $g|D{Q}^{\ast}$, it follows that $g=\operatorname{rgcd}(g,D)\operatorname{lgcd}(g,{Q}^{\ast})=\operatorname{rgcd}(g,D_{2})\operatorname{lgcd}(g,{Q}^{\ast})$ by Lemma 2, thus $g|D_{2}{Q}^{\ast}$. Therefore, $g^{2}$ divides $Q{D}^{\ast}Q_{L}$ and ${Q}^{\ast}_{L}{D}{Q}^{\ast}$. Therefore, $c$ divides $D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}$ because we have $D^{\prime}_{R}{D^{\prime}_{R}}^{\ast}={D}^{\ast}_{L}Q{Q}^{\ast}D_{L}+Q_{L}{Q}^{\ast}_{L}\frac{D{D}^{\ast}}{g^{2}}+\frac{c}{g^{2}}\left({D}^{\ast}_{L}Q{D}^{\ast}Q_{L}+{Q}^{\ast}_{L}{D}{Q}^{\ast}D_{L}\right)$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n\in\mathbb{N}$ with $\operatorname{mrpf}(Q)=1$ and $\nu(M)=c^{2}\nu(Q)=N^{n}$ for some irreducible, quadratic polynomial $N\in\mathbb{R}[t]$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$. If $cg$ divides $\nu(D)$, then $M$ admits a factorization into monic linear factors.